OpenStudy (anonymous):

Anthony leaves his house on his bike traveling at 18 mph. Thirty minutes later, his brother Chandler takes after him at 24 mph. How many hours will it take Chandler to catch up to Anthony?

OpenStudy (anonymous):

I've got d=18*30+24*30 540+720 and my answer I got was 1260

OpenStudy (anonymous):

You are thinking about it the wrong way. You want the distances to be equal, not the times. some distance d = the time they have been traveling times the speed they were traveling at. so for Anthony the equation is d=(30 + t) * 18 because he has been traveling for 30 mins longer than Chandler for Chandler it is d=t*24 The distances are the same so you can set 24t=18(30+t) and solve that t= 90 mins for Chandler to catch Anthony

OpenStudy (anonymous):

Thank you.!